### v3

@@ Methods
All methods were run while turning off the display of visualizations on the screen and without saving the visualizations to the disk. To do so, some of the code of Detectron2 and the MMPose environment was modified, as no parameter was provided to turn them off. Further modifications were made to the Detectron2 code to output the keypoint coordinates as a file, as it did not provide a way to get such output by default. The ”lap” and SyRIP datasets contain non-sequential images extracted from different videos of different infants at different ages. The video input option was thus not available. On VITPose which in the version of MMPose we used could only process videos, we modified the code of MMPose’s ViTPose to accept images as inputs, without modifying any of the underlying processing done by ViTPose. While that enabled us to present ViTPose results on these datasets, ViTPose might internally only expect videos and have processing steps that enhance estimations assuming a time coherence between two processed frames. As there was not any parameter available to indicate to ViTPose a lack of sequential coherence, we avoided using the image processing steps for the other two datasets that had videos that could be processed (”supine”, and MINI-RGBD). The performance of ViTPose on the ”lap” and SyRIP dataset could be lower than would be in a real application scenario with full video recordings to process. All modifications are effective in the Dockers that we share (at https://hub.docker.com/u/humanoidsctu) and are indicated in the documentation.
@@ Results
The results for both real and synthetic infants are shown for each metric described above. In each subsection, we provide the results for the manually-annotated dataset of real infants first: in order, our ”supine” and ”lap” datasets, then the SyRIP dataset, and then the results for the MINI-RGBD dataset of synthetic infants are presented last. Due to the varying complexity of infant postures across the datasets, we provide detailed results for individual sequences in the Supplementary Materials when relevant.
@@
The average OKS values for each dataset (average across individual infant images) are shown in Tab. 2 (see Supplementary Tables ST. 5 and ST. 6 to see the full details per video on the ”supine” and MINI-RGBD datasets).
For real infant images in ”supine” position, ViTPose and HRNet TD have the highest OKS, followed by HRNet BU. The results between frame-by-frame and video inputs are almost identical, except for OpenPose, for which image input leads to better results. On the ”lap” and SYRIP datasets, we observe a significant drop of performance for all methods, with particularly high variance, and wider gaps of performance between the methods. HRNet BU seems to perform the best on the ”lap” dataset, while ViTPose leads on the SyRIP dataset. For synthetic infants, VitPose and HRnet TD also show the best results. The differences between image and video inputs are narrow but consistent between the different methods, with a slight edge for image inputs. Synthetic infants in supine position are more difficult to process for all methods, in particular for HRNet BU, compared to the real infants in supine position, with the exceptions being DeepLabCut and MediaPipe, which perform better on synthetic infants, but the performance is overall higher than for the ”lap” or the SyRIP datasets.
@@
Synth. Images 0.84$\pm$0.11 N/A 0.84$\pm$0.10 0.50$\pm$0.21 0.83$\pm$0.16 0.88$\pm$0.7 0.83$\pm$0.12 N/A
MINI-RGBD Videos 0.81$\pm$0.15 0.43$\pm$0.20 0.81$\pm$0.12 0.47$\pm$0.22 0.81$\pm$0.17 0.86$\pm$0.10 0.81$\pm$0.12 0.87$\pm$0.07
Table 2: Average OKS values across individual images for each method, dataset and input type.
@@
For real infants, ViTPose has the highest AP and AR. Considering only AP, ViTPose is followed by HRNet BU and then by OpenPose. Lower AP values for HRNet TD and Detectron2 in our supine dataset are influenced by a large number of redundant high-confidence detections (see Tab. 6). Considering only AR, ViTPose is followed by HRNet TD and HRNet BU. Whether the inputs are images (frames) or videos, the results are roughly similar except for OpenPose, though it seems that, generally, image inputs lead to better results (except for HRNet TD, for which video input seems preferred). There is a large drop of AP and AR in the ”Lap” and SyRIP datasets compared to the supine infants.
For synthetic infants, both the AP and AR values are lower than those in the supine dataset of real infants; it seems that the methods have more difficulty processing synthetic infants, except for DeepLabCut and MediaPipe. Considering only AP, ViTPose is the best method, followed by OpenPose. Considering only AR, HRNet TD is the best method, followed by ViTPose. We observe again that despite the lower scores, the synthetic infants of the MINI-RGBD datasets, which are also in supine position and seen from above, are better estimated by all methods than the real infants in the lap and in the SyRIP datasets.
@@
Neck-MidHip errors are shown in Fig. 2. DeepLabCut errors are not shown, as they create large radius circles that impede readability of the figures. The version of the figures including DeepLabCut is available in the Supplementary Figure SF. 2. Across all methods, the estimations of the eyes and nose are the most accurate, leading to the smallest errors, followed by shoulders and wrists. The positions of the hips and knees are detected with the highest errors. We can observe the increase in errors between the real infants in the ”supine” condition, the MINI-RGBD dataset, the SyRIP dataset, and finally the ”Lap” dataset where the errors make the image difficult to read. The details of the overall average Neck-MidHip errors across all keypoints with their standard deviations for each method are shown in Tab.  4. The variability of the errors is high, often above one third, and sometimes even half, of the average error. This might be explained by the large variability between each keypoint type as can be seen on Fig. 2, or Supplementary Materials Tab. ST 7, 8, 9 and 10.
Overall, for real infants, the best method is ViTPose, followed by HRNet TD. However, in the ”Lap” dataset, while the performance of all the methods drops significantly, ViTPose and the other top-down methods seem to struggle particularly compared to their performance on the other datasets. HRNet BU has the lowest errors among all methods, though the errors are high in absolute terms: an average error of 24% of the Neck-MidHip length. A less dramatic but general drop of performance is observed in the SyRIP dataset as well, where the best performing method is ViTPose with relatively high average errors at 12% of the Neck-MidHip length. For synthetic infants, the best method is HRnet TD, followed by ViTPose, although HRNet TD has slightly higher standard deviations.
Figure 2: Neck-midhip errors for each keypoint with available ground truth. The centre of the circles is the ground-truth position for that keypoint. The radius of each circle shows the average error as a proportion of the Neck-MidHip length (see Methods). The colors represent separately each pose estimation method. a) Real Infants ”supine”, image input; b) Real infants ”supine”, video input; c) MINI-RGBD synthetic infants, image input; d) MINI-RGBD synthetic infants, video input; d) Real infants ”lap”; and e) Real infants, SyRIP. DeepLabCut estimations leading to much bigger errors not shown (see Supplementary Fig. SF. 2).
@@
For real infants, the methods that never miss detections or keypoints are DeepLabCut and Detectron2, while HRNet BU is very close behind, with only a few misses that are lost when rounded at less than 2 decimal points. ViTPose and HRNet TD miss an amount of data that could be considered acceptable, below 0.5%. On the other hand, OpenPose, MediaPipe, and AlphaPose display high amounts of missing data. We observe a tendency to miss smaller amounts of data with video input compared to image input. OpenPose reaches more than 50%, the highest number of missing data overall, on the Lap dataset.
For synthetic infants, the methods missing data are OpenPose, MediaPipe, and AlphaPose, though to a lesser degree than for real infants. Contrary to the real infants, we observe a tendency to miss smaller amounts of data with image input compared to video input.
@@
The percentages of redundant detections for each method are shown in Tab. 6. The details for each individual video can be found in the Supplementary Tables ST. 11 and 12. DeepLabCut, MediaPipe and OpenPose cannot have redundant detections by design, though for the case of the ”Lap” dataset, they could have minus detections. For the other methods, on the ”supine” real infants dataset, the methods with the least amount of redundant data are AlphaPose and HRNet BU. Except for HRNet TD, we observe a tendency, especially for Detectron2, to produce more redundant detections with video input. On the ”Lap” dataset, HRNet BU, Detectron2, and HRNet TD display higher rates of redundant detections, while on the contrary, MediaPipe, ViTPose, and AlphaPose seem to not always detect both individuals and only detect one of them (often the adult) for high percentages of the images. OpenPose is the method that overall has the least difference in detections compared to the expected amount. On SyRIP, all methods except MediaPipe and OpenPose display high amounts of redundant detections.
@@
Overall, Detectron2 and HRNet TD overall provide disproportionate amounts of redundant detections, especially on real infants, though HRNet BU has the most redundant detections in the ”Lap” context, where, on the contrary, MediaPipe struggles the most at providing at least 2 detections (one of each individual on the image).
In the ”supine” and MINI-RGBD datasets, it is possible to look at the details per video (see Supplementary Materials Tables ST 11 and 12.) In doing so, we see that the distribution of redundant detections is not uniform. Detectron2, HRNet TD and ViTPose show higher redundant detections for synthetic infants that are considered more difficult in the MINI-RGBD dataset, while HRNet BU shows higher redundant detections for the easier synthetic infants.
@@
For real infants, in the ”supine” context, the method with the highest Correlation Coefficients is HRNet BU for both input types. With image input, it is closely followed by OpenPose. The correlations are in the correct direction for all methods. It is interesting to note that on the ”Lap” dataset, some methods maintain correlation scores in line with the other datasets, while others have lower correlations (AlphaPose, OpenPose, and ViTPose). In contrast, on the SyRIP dataset, most methods display higher correlations, especially AlphaPose. Still, HRNet BU is the method with the highest correlations in all the real infants datasets. For synthetic infants, the methods with the highest Correlation Coefficients is OpenPose for image inputs. For video inputs, HRNet BU has the highest correlations, closely followed by OpenPose.
@@
Table 7: Spearman correlations between individual frames OKS values and score values for each method, dataset and input type. All p-values $<$ 0.005 except for DeepLabCut on real infant videos with video input.
@@ Conclusion
This article provided an empirical comparison of seven state-of-the-art human pose estimation methods—trained on images of adults during various activities, typically in upright positions—on several datasets of infant images, including infants under 7 months of age in supine position, infants between 17 weeks to 18 months in close proximity with a parent, and infants below 1 year old in a variety of contexts and postures. Our results are summarized below.
@@
First, we conclude that state-of-the-art human pose estimation methods work well to estimate infant poses without the need for additional training or fine-tuning for infants in ”supine” position. An overview of the estimation accuracy for individual keypoints on the body is provided in Fig. 2, and in Supplementary Materials Tables ST 6-9. Overall accuracy estimations are available in Tables  2,  3 using standard performance metrics (OKS, AP/AR) and in Table 4 using an infant-size centred metric. ViTPose has the best accuracy [49], followed by HRNet [47, 48] (top-down variant). The other networks tested (HRNet bottom-up, AlphaPose [31], Detectron2 [44], and OpenPose [28]) have higher errors. The DeepLabCut environment [51], with human pose estimation using DeeperCut [43], as well as MediaPipe [45] with BlazePose [46] does not provide competitive results at all. It is coherent that ViTPose, being based on one of the newest state-of-the-art neural network architecture, Transformers, which introduce an ”attention” mechanism, as well as being trained on more datasets that the other methods, ended up with the best performance. These results are in line with [42] who used the Percentage of Correct Keypoints (PCK), an accuracy-based evaluation that can be compared to our Neck-MidHip errors, as they also base it on the infant’s approximate torso length. The performance rank order of the methods we commonly compare are the same, with the identical conclusion that ViTPose performs the best. The error levels also converge towards similar levels: they find that 84.6% of the ViTPose estimations are equal or below 10% of the infant torso length, while we find that 68% of the ViTPose estimations are within 3.3% to 8.7% of the infant torso length. However, compared to [42], who only used a single accuracy-based evaluation, PCK, we complemented our results using several accuracy-based metrics, including the standard AP and AR, with other analyses such as the amount of missing data or of produced redundant detection, the relation between scores and confidence values with accuracy-based performance, as well as the processing speed of each method. We also included some methods that we observed to be often used by infant development researchers (AlphaPose and DeepLabCut in particular).
@@
On the more complex ”Lap” and SyRIP datasets, ViTPose and either OpenPose or HRNet depending on the metric and dataset have the highest performance. However, for all methods, the drop in performance is drastic, for example losing between 32 to 65 points in Average Precision and no method reaching errors lower than an average of 12% of the infants’ torso length.
@@
Second, we complemented accuracy-based performance analysis with additional criteria that are important for the practical applications of the methods. The accuracy (OKS, Neck-midhip errors) cannot be computed if keypoints are not detected and does not penalize the case where multiple people are erroneously detected (though AP partially consider these cases). Moreover, both missing and redundant detections negatively impact any downstream motion analysis. Detectron2 and HRNet (top-down) are in particular susceptible to redundant detections. OpenPose, MediaPipe and AlphaPose have high rates of missing data (missed detections of the whole infant or of individual keypoints). In the case of infant-parent (dyadic) interaction settings, such as in our ”Lap” dataset, disentangling individual detections is not trivial, even when using dual RGB-D camera setup in a laboratory setting [59]. One reason for this is that the methods are biased towards adults instead of infants, and will sometimes only detect the parent. This is particularly noticeable when looking at methods that are unable to provide as many detections as expected in this context; see Table  6.
@@
It is also important to consider both accuracy-based metrics (AP/AR, Neck-MidHip, OSK) together with higher level metrics (missing data, redundant detections): for example, OpenPose, MediaPipe and AlphaPose have high amounts of missing data. Missing keypoints, in particular, which happen for bottom-up methods either because the methods cannot detect them in the image, or because their tentative placement was too far from the model’s priors and were ignored, can not be evaluated by accuracy-based errors. In contrast, top-down methods would have a tendency to provide all keypoints even when misplaced, giving a lower confidence or score value to flag this. While confidence thresholds can be arbitrarily chosen to cull and remove those keypoints, which some methods do internally, we decided to not set such an arbitrary threshold and to not remove them in our results for the methods that did not, making these keypoints impact the accuracy-based metrics instead of showing as missing data. The most striking example might be OpenPose, a Bottom-Up method, which misses up to 54% of the data, mainly as missing keypoints, in the ”Lap” dataset, yet manages to reach the second highest AP values and the second lowest average Neck-MidHip errors for this dataset.
@@
Researchers in behavioral science are not experts in machine learning and computer vision and cannot keep up with the rapid progress in human pose estimation methods. Hence, they choose their tools based on methods that have been used in the past (in the context of infants, OpenPose was used in [29, 33, 30]; AlphaPose in [20]). The DeepLabCut environment [51] is the frequent tool of choice in behavioral science. Here we show that for the estimation of the pose of infants in supine position, DeepLabCut with human pose estimation using DeeperCut [43] currently does not provide satisfactory performance at all. OpenPose and AlphaPose may be employed but fall behind ViTPose and HRNet (top-down) in accuracy, as well as missed data. We recommend the community to use the currently best-performing tools and profit from the Docker containers we released.

### v4

@@ Methods
Combined Performance Evaluation (CPE)
It is desirable to have a metric that combines the metrics above. However, several metrics (OKS, AP and AR, Neck-MidHip error) are concerned with the accuracy of the keypoint detections in the image and hence are not independent. Additional metrics we consider are supplementary and their relevance depends on the context (redundant detections, confidence-OKS correlations, processing speed). It would still be useful to combine one of the accuracy metrics with the the percentage of missing data.
AP and AR could be combined into a single metric like the F1 score, but cannot be easily combined with the percentage of missing data as AR already considers missing detections. OKS uses weighted coefficients for each keypoint type and has a non-linear relationship with the absolute errors. As such, we preferred to opt for the metric that had the least overlap with others and was the most neutral: Neck-MidHip errors.
Therefore, we propose the Combined Performance Evaluation (CPE) as a metric combining the accuracy of detections based on the Neck-MidHip errors with the percentage of missing data through Equation 4.
$f\colon x\rightarrow(1-MIN(1,(x/(c\times 100))))\in[0,1]\qquad\text{CPE}_{m}=(f(nmh\_error_{m})+f(mdata_{m}))/2$ (4)
where $f(x)$ is a function that turns the positive percentage-based values of the Neck-MidHip errors and missing data into a normalized scale between 0 and 1. Here, 1 corresponds to perfect estimations with no missing data and no estimation error. The scaling coefficient $c$ determines the limit (threshold) value that allows non-zero output for any $x$ value below the threshold. Here, we set $c=0.5$ so that the threshold (divisor) is $50$ by considering that estimation errors above 50% of the torso length and missing data on more than 50% of a recording are not acceptable for the use of pose estimation methods for further analysis. This coefficient penalizes larger errors while providing finer granularity of the final values to separate methods that have similar performance. The final score is computed as the mean between the scaled and normalized values obtained for the Neck-MidHip errors and the missing data.
@@
All methods were run while turning off the display of visualizations on the screen and without saving the visualizations to the disk. To do so, some of the code of Detectron2 and the MMPose environment was modified, as no parameter was provided to turn them off. Further modifications were made to the Detectron2 code to output the keypoint coordinates as a file, as it did not provide a way to get such output by default.
The “Lap” and SyRIP datasets contain non-sequential images extracted from different videos of different infants at different ages. The video input option was thus not appropriate. On ViTPose, which in the version of MMPose we used could only process videos, we modified the code of MMPose’s ViTPose to accept images as inputs, without modifying any of the underlying processing done by ViTPose. While that enabled us to present ViTPose results on these datasets, an important comparison due to ViTPose being the contender state-of-the-art method in adult pose estimation, ViTPose might internally only expect videos and have processing steps that enhance estimations assuming a time coherence between two processed frames. As there was no parameter available to indicate to ViTPose a lack of sequential coherence, we avoided using the image processing steps for the other two datasets that had videos that could be processed (“Supine”, and MINI-RGBD). The performance of ViTPose on the “Lap” and SyRIP dataset could be lower than in a real application scenario with full video recordings to process. All modifications are effective in the Dockers that we share (at https://hub.docker.com/u/humanoidsctu) and are indicated in the documentation.
@@ Results
The results for both real and synthetic infants are shown for each metric described above. In each subsection, we provide the results for the manually-annotated dataset of real infants first: in order, our “Supine” and “Lap” datasets, then the SyRIP dataset, and then the results for the MINI-RGBD dataset of synthetic infants are presented last. Due to the varying complexity of infant postures across the datasets, we provide detailed results for individual sequences in the Supplementary Materials when relevant.
@@
The average OKS values for each dataset (average across individual infant images) are shown in Tab. 2 (see Supplementary Tables ST. 6 and ST. 7 to see the full details per video on the “Supine” and MINI-RGBD datasets).
For real infant images in supine position, ViTPose and HRNet TD have the highest OKS, followed by HRNet BU. The results between frame-by-frame and video inputs are almost identical, except for OpenPose, for which image input leads to better results. On the “Lap” and SyRIP datasets, we observe a significant drop of performance for all methods, with particularly high variance, and wider gaps of performance between the methods. HRNet BU seems to perform the best on the “Lap” dataset, while ViTPose leads on the SyRIP dataset. For synthetic infants, ViTPose and HRNet TD also show the best results. The differences between image and video inputs are narrow but consistent between the different methods, with a slight edge for image inputs. Synthetic infants in supine position are more difficult to process for all methods, in particular for HRNet BU, compared to the real infants in supine position, with the exceptions being DeepLabCut and MediaPipe, which perform better on synthetic infants, but the performance is overall higher than for the “Lap” or the SyRIP datasets.
@@
[…]
Table 2: Average OKS values across individual images for each method, dataset and input type. N/A values correspond to methods not evaluated on image inputs as they only accept video input (see Sec. “Comparison metrics” first paragraph and Sec. “Hardware and methods’ code modifications” for more details)
@@
For real infants, ViTPose has the highest AP and AR. Considering only AP, ViTPose is followed by HRNet BU and then by OpenPose. Lower AP values for HRNet TD and Detectron2 in our “Supine” dataset are influenced by a large number of redundant high-confidence detections (see Tab. 6). Considering only AR, ViTPose is followed by HRNet TD and HRNet BU. Whether the inputs are images (frames) or videos, the results are roughly similar except for OpenPose, though it seems that, generally, image inputs lead to better results (except for HRNet TD, for which video input seems preferred). There is a large drop of AP and AR in the “Lap” and SyRIP datasets compared to the supine infants.
For synthetic infants, both the AP and AR values are lower than those in the “Supine” dataset of real infants; it seems that the methods have more difficulty processing synthetic infants, except for DeepLabCut and MediaPipe. Considering only AP, ViTPose is the best method, followed by OpenPose. Considering only AR, HRNet TD is the best method, followed by ViTPose. We observe again that despite the lower scores, the synthetic infants of the MINI-RGBD datasets, which are also in supine position and seen from above, are better estimated by all methods than the real infants in the “Lap” and in the SyRIP datasets.
@@
Neck-MidHip errors are shown in Fig. 2. DeepLabCut errors are not shown, as they create large radius circles that impede readability of the figures. The version of the figures including DeepLabCut is available in the Supplementary Materials Figure SF. 2. Across all methods, the estimates of the eyes and nose are the most accurate, leading to the smallest errors, followed by shoulders and wrists. The positions of the hips and knees are detected with the highest errors. We can observe the increase in errors between the real infants in the “Supine” dataset, the MINI-RGBD dataset, the SyRIP dataset, and finally the “Lap” dataset where the errors make the image difficult to read. The details of the overall average Neck-MidHip errors across all keypoints with their standard deviations for each method are shown in Tab.  4. The variability of the errors is high, often above one third, and sometimes even half, of the average error. This might be explained by the large variability between each keypoint type as can be seen on Fig. 2, or Supplementary Materials Tables ST 8-11.
Overall, for real infants, the best method is ViTPose, followed by HRNet TD. However, in the “Lap” dataset, while the performance of all the methods drops significantly, ViTPose and the other top-down methods seem to struggle particularly compared to their performance on the other datasets. HRNet BU has the lowest errors among all methods, though the errors are high in absolute terms: an average error of 24% of the Neck-MidHip length. A less dramatic but general drop of performance is observed in the SyRIP dataset as well, where the best performing method is ViTPose with relatively high average errors at 12% of the Neck-MidHip length. For synthetic infants, the best method is HRnet TD, followed by ViTPose, although HRNet TD has slightly higher standard deviations.
As an additional comparison, we tested to what extent a “mixture of experts” approach—combining several of the estimation methods—could yield better performance. We tested averaging the estimates from the best-performing methods: ViTPose, HRNet TD and HRNet BU. The details can be found in the Supplementary Materials, sections “Mixture of experts estimates” and “Results summary for the ‘mixture of experts’ estimates”. Overall, the averaged estimates are worse than the best individual method. The mixture could make the estimates more robust in face of missing data though.
[…]
Figure 2: Neck-midhip errors for each keypoint with available ground truth. The centre of the circles is the ground-truth position for that keypoint. The radius of each circle shows the average error as a proportion of the Neck-MidHip length (see Methods). The colors represent separately each pose estimation method. a) Real Infants “Supine”, image input; b) Real infants “Supine”, video input; c) MINI-RGBD synthetic infants, image input; d) MINI-RGBD synthetic infants, video input; e) Real infants “Lap”; and f) Real infants, SyRIP. DeepLabCut estimates leading to much higher errors not shown (see Supplementary Fig. SF. 2).
@@
For real infants, the methods that never miss detections or keypoints are DeepLabCut and Detectron2, while HRNet BU is very close behind, with only a few misses that are lost when rounded at less than 2 decimal points. ViTPose and HRNet TD miss an amount of data that could be considered acceptable, below 0.5%. On the other hand, OpenPose, MediaPipe, and AlphaPose display high amounts of missing data. We observe a tendency to miss smaller amounts of data with video input compared to image input. OpenPose reaches more than 50%, the highest number of missing data overall, on the “Lap” dataset.
For synthetic infants, the methods that produce missing data are OpenPose, MediaPipe, and AlphaPose, though to a lesser degree than for real infants. Contrary to the real infants, we observe a tendency to miss smaller amounts of data with image input compared to video input.
@@
The percentages of redundant detections for each method are shown in Tab. 6. The details for each individual video can be found in the Supplementary Tables ST. 12 and 13.
DeepLabCut, MediaPipe and OpenPose cannot have redundant detections by design, though for the case of the “Lap” dataset, they could have minus detections. For the other methods, on the “Supine” real infants dataset, the methods with the least amount of redundant data are AlphaPose and HRNet BU. Except for HRNet TD, we observe a tendency, especially for Detectron2, to produce more redundant detections with video input.
On the “Lap” dataset, HRNet BU, Detectron2, and HRNet TD display higher rates of redundant detections, while on the contrary, MediaPipe, ViTPose, and AlphaPose seem to not always detect both individuals and only detect one of them (often the adult) for high percentages of the images. OpenPose is the method that overall has the least difference in detections compared to the expected amount. On SyRIP, all methods except MediaPipe and OpenPose display high amounts of redundant detections.
@@
Overall, Detectron2 and HRNet TD overall provide disproportionate amounts of redundant detections, especially on real infants, though HRNet BU has the most redundant detections in the “Lap” context, where, on the contrary, MediaPipe struggles the most at providing at least 2 detections (one of each individual on the image).
In the “Supine” and MINI-RGBD datasets, it is possible to look at the details per video (see Supplementary Materials Tables ST 11 and 12.) In doing so, we see that the distribution of redundant detections is not uniform. Detectron2, HRNet TD and ViTPose show higher redundant detections for synthetic infants that are considered more difficult in the MINI-RGBD dataset, while HRNet BU shows higher redundant detections for the easier synthetic infants.
@@
For real infants, in the “Supine” context, the method with the highest Correlation Coefficients is HRNet BU for both input types. With image input, it is closely followed by OpenPose. The correlations are in the correct direction for all methods. It is interesting to note that on the “Lap” dataset, some methods maintain correlation scores in line with the other datasets, while others have lower correlations (AlphaPose, OpenPose, and ViTPose). In contrast, on the SyRIP dataset, most methods display higher correlations, especially AlphaPose. Still, HRNet BU is the method with the highest correlations in all the real infants datasets.
For synthetic infants, the methods with the highest Correlation Coefficients is OpenPose for image inputs. For video inputs, HRNet BU has the highest correlations, closely followed by OpenPose.
@@
Table 7: Spearman correlations between individual frames OKS values and score values for each method, dataset and input type. All p-values << 0.005 except for DeepLabCut on real infant videos with video input.
@@
Combined Performance Evaluation (CPE)
The CPE values, combining Neck-MidHip errors (error-based metric) and the percentage of missing data, are shown in Tab. 9.
On real infants, ViTPose is the best performing method overall due to its lower estimation errors and a low amount of missing data, except on the “Lap” dataset where it has higher errors and HRNet BU leads with the highest CPE value. The differences appear for other methods that are closer on their estimation errors but vary more largely on their amounts of missing data. The performance gap between the methods shows more clearly when looking at the more challenging SyRIP and “Lap” datasets, with the extreme example of OpenPose that had the second lowest Neck-MidHip errors on the “Lap” dataset, right after HRNet BU, but ends up with the lowest CPE value among all methods on this dataset, as it misses 54% of the data, while HRNet BU does not miss any data and has the highest CPE value.
For synthetic infants, the CPE values are overall very close to the real infants in supine position, with ViTPose and HRNet TD coming out on top.
Real Images 0.84 N/A 0.90 0.53 0.92 0.93 0.81 N/A
Supine Videos 0.85 0.50 0.90 0.55 0.92 0.90 0.79 0.94
Lap Images 0.61 N/A 0.58 0.44 0.76 0.57 0.25 0.6
SyRIP Images 0.82 N/A 0.81 0.30 0.84 0.86 0.71 0.88
Synth. Images 0.85 N/A 0.90 0.63 0.88 0.92 0.86 N/A
MINI-RGBD Videos 0.79 0.57 0.89 0.60 0.87 0.91 0.85 0.92
Table 9: Combined Performance Evaluation values based on Neck-MidHip errors and Missing Data percentages for each method, dataset and input type.
@@ Conclusion
This article provided an empirical comparison of seven state-of-the-art human pose estimation methods—trained on images of adults during various activities, typically in upright positions—on several datasets of infant images, including infants under 7 months of age in supine position, infants between 17 weeks to 18 months sitting and in close proximity with a parent, and infants below 1 year old in a variety of contexts and postures. Our results are summarized below.
@@
First, we conclude that state-of-the-art human pose estimation methods work well to estimate infant poses even without additional training or fine-tuning for infants in supine position. An overview of the estimation accuracy for individual keypoints on the body is provided in Fig. 2, and in Supplementary Materials Tables ST 6-9. Overall accuracy evaluations are available in Tables 2 and 3 using standard performance metrics (OKS, AP/AR) and in Tab. 4 using an infant-size centred metric. ViTPose has the best accuracy [49], followed by HRNet [47, 48] (top-down variant). The other networks tested (HRNet bottom-up, AlphaPose [31], Detectron2 [44], and OpenPose [28]) have higher errors. The DeepLabCut environment [51], with human pose estimation using DeeperCut [43], as well as MediaPipe [45] with BlazePose [46] do not provide competitive results at all. It is coherent that ViTPose, being based on one of the newest state-of-the-art neural network architecture, Transformers, which introduces an ”attention” mechanism, as well as being trained on more datasets than the other methods, ended up with the best performance.
These results are in line with [42] who used Percentage of Correct Keypoints (PCK), an accuracy-based evaluation that can be compared to our Neck-MidHip errors to some extent, as they also base it on the infant’s approximate torso length. The performance rank order of the methods we commonly compare are the same, with the identical conclusion that ViTPose performs the best. The error levels also converge towards similar levels: they find that 84.6% of the ViTPose estimates are equal or below 10% of the infant torso length, while we find that 68% of the ViTPose estimates are within 3.3% to 8.7% of the infant torso length. However, compared to [42], who only used this single accuracy-based evaluation, PCK, we complemented our results using several accuracy-based metrics, including the standard AP and AR, with other analyses such as the amount of missing data or of produced redundant detection, the relation between scores and confidence values with accuracy-based performance, as well as the processing speed of each method. We also included the methods that we observed to be often used by infant development researchers (AlphaPose and DeepLabCut in particular, in addition to OpenPose).
@@
On the more complex “Lap” and SyRIP datasets, ViTPose and either OpenPose or HRNet depending on the metric and dataset have the highest performance. However, for all methods, the drop in performance is drastic, for example losing between 32 to 65 points in Average Precision and no method reaching errors lower than an average of 12% of the infants’ torso length.
@@
Second, we complemented accuracy-based performance analysis with additional criteria that are important for the practical applications of the methods. The accuracy (OKS, Neck-midhip errors) cannot be computed if keypoints are not detected and does not penalize the case where multiple people are erroneously detected (though AP partially consider these cases). Moreover, both missing and redundant detections negatively impact any downstream motion analysis. Detectron2 and HRNet (top-down) are in particular susceptible to redundant detections. OpenPose, MediaPipe and AlphaPose have high rates of missing data (missing detections of the whole infant or of individual keypoints). In the case of infant-parent (dyadic) interaction settings, such as in our “Lap” dataset, disentangling individual detections is not trivial, even when using dual RGB-D camera setup in a laboratory setting [60]. One reason for this is that the methods are biased towards adults instead of infants, and will sometimes only detect the parent. This is particularly noticeable when looking at methods that are unable to provide as many detections as expected in this context; see Tab. 6.
@@
It is also important to consider both accuracy-based metrics (AP/AR, Neck-MidHip, OSK) together with higher level metrics (missing data, redundant detections): for example, OpenPose, MediaPipe and AlphaPose have high amounts of missing data. Missing keypoints, in particular, which happen for bottom-up methods either because the methods cannot detect them in the image, or because their tentative placement was too far from the model’s priors and were ignored, can not be evaluated by accuracy-based errors. The most striking example might be OpenPose, a Bottom-Up method, which misses up to 54% of the data, mainly as missing keypoints, in the “Lap” dataset, yet manages to reach the second highest AP values and the second lowest average Neck-MidHip errors for this dataset. In contrast, top-down methods would have a tendency to provide all keypoints even when misplaced, giving a lower confidence or score value to flag this. While confidence thresholds can be arbitrarily chosen to remove those keypoints, which some methods do internally, we decided to not set such an arbitrary threshold and to not remove them in our results for the methods that did not do so by default internally, making these keypoints impact the accuracy-based metrics instead of showing as missing data in our results.
For this purpose, we introduced a Combined Performance Evaluation (CPE) that merges accuracy-based Neck-MidHip errors with the percentage of missing data into a single value. OpenPose drops to being the worst performing method on the “Lap” dataset in this Combined Performance Evaluation. Using this evaluation, we find again that ViTPose displays the best performance except for the “Lap” dataset, most likely due to a strong bias towards the detection of the adult parent.
@@
Researchers in behavioral science are not experts in machine learning and computer vision and cannot keep up with the rapid progress in human pose estimation methods. Hence, they choose their tools based on methods that have been used in the past (in the context of infants, OpenPose was used in [33, 30, 29]; AlphaPose in [20]). The DeepLabCut environment [51] is a frequent tool of choice in behavioral science. Here we show that for the estimation of the pose of infants in supine position, DeepLabCut with human pose estimation using DeeperCut [43] currently does not provide satisfactory performance at all. OpenPose and AlphaPose may be employed but fall behind ViTPose and HRNet (top-down) in accuracy, as well as percentage of missing detections of infants or body keypoints.
While fine-tuning these methods with infant data can improve their performance [42, 40, 29], due to the limited amount of annotated infant data, it seems to easily and quickly overfit the models to the specificities of the training subset and to reduce the performance when used on other image sets [42]. The amount of “difference” that leads to under performance after fine-tuning is difficult to quantify due to the nature of these models (“black-box” deep learning networks), and depends on the quantity of available data and how the training dataset is separated from the rest of the data on which the model will be applied to. We recommend the community to use the currently best-performing tools and profit from the Docker containers we released.